# Batch Estimation of a Steady, Uniform, Flow-Field from Ground Velocity and Heading Measurements

### v1

@@ 4 Comparison using Simulated Data
To compare the performance of the four methods described in Secs. 3 and 2, a Monte Carlo experiment was conducted using simulated data. The data was simulated for three different noise settings, $\sigma=\{0.01,0.05,0.10\}$ m/s, and three different heading angle changes, $\Delta\psi=\{2\pi,3\pi/2,\pi\}$. For each combination of $\sigma$ and $\Delta\psi$ a total of 250 unique datasets of $N=100$ noisy samples were generated and transformed to the appropriate form for each algorithm. Each dataset was generated by randomizing the true parameters $(v,w,\theta)$ and drawing them uniformly at random from the interior of the polytope represented by the constraints with $v_{\rm min}=0.5$ and $v_{\rm max}=5$ m/s. For each dataset the circular fit algorithm, optimization with $(\dot{x},\dot{y})$ data, and optimization with $(\dot{v}_{g},\psi)$ data were tested. The quadratic fit approach was evaluated only for the $\Delta\psi=2\pi$ case. The estimate each algorithm produced was compared to the actual value and the error was recorded. The mean error, average over 250 trials, is shown in Fig. 6.
As expected, the error for each algorithm increased as the measurement noise is increased and as the range of data is reduced. The quadratic fit method produced the largest errors (e.g., up to 18 deg in mean current direction error for the highest noise setting). The optimization with the $(v_{g},\psi)$ data was not as accurate as the optimization with $(\dot{x},\dot{y},\psi)$ data. This is not surprising since more information is available when the components of the ground velocity are known, rather than the magnitude. The circle fit, which uses $(\dot{x},\dot{y})$ data, in most cases had an intermediate error in comparison to the the two optimization-based methods. The optimization method using $(\dot{x},\dot{y},\psi)$ data produced the most accurate results — for example, in the case of data collected for $\Delta\psi=2\pi$ maneuvers the vehicle and current speed errors were $<$ 0.01 m/s and the current direction error was $<$ 5 degrees.
@@ 5 Comparison using Experimental Data
The four methods described in Secs. 3 and 2 were also evaluated using experimental data obtained by a Bluefin-21 unmanned underwater vehicle operated by the U.S. Naval Research Laboratory. The Bluefin-21 is a 20 ft length, 21 inch diameter vehicle developed by Bluefin Robotics that features a fiber-optic gyroscope inertial navigation system and Doppler velocity log (DVL) navigation suite. The vehicle was deployed on June 22, 2016 near Boston Harbor southeast of Nahant Bay. The vehicle was programmed to execute a series of circular orbits around points of interest to test an onboard sonar system. The ground track of the vehicle is shown in Fig. 7 along with the flow direction estimated by the optimization method with $(\dot{x},\dot{y},\psi)$ data.
@@
The radii of the smaller and larger orbits were 70 and 120 meters, respectively. The data under analysis was truncated to be within 10 meters of these orbits and input into all four estimation algorithms. An example of the dataset along one such orbit is shown in Fig. 8. The parameters estimated by each algorithm were used to superimpose a $v_{g}(\psi)$ curve over the data points. The dataset exhibits outliers that do not conform to the model (e.g., due to transient motions as the vehicle enters and exits each orbit). To address this issue the optimization algorithms are wrapped in a random sample consensus (RANSAC) algorithm outer loop.
The estimated current direction and magnitude was also compared to historical data from an instrumented buoy (identifier: BOS1132) deployed by the National Oceanic and Atmospheric Administration (NOAA) located in Stellwagen Bank about 51 km away directly east from the testing area (15nm NNE of Race Point at a depth of 27ft). The results show good agreement with the ebb and flow tidal directions reported for the buoy. The current magnitude matches the general trend and variations may be due to the spatial separation of the buoy and vehicle leading to different tidal flows closer to shore.
Figure 8: Example experimental data along a single orbit with overlayed $v_{g}(\psi)$ curves drawn with parameters from the four estimation methods. Refer to Fig. 6 for a legend.
Figure 9: Comparison of estimation methods with NOAA buoy data. Refer to Fig. 6 for a legend.
@@ 6 Conclusion
Three batch estimation methods were presented that determine the direction and magnitude of a steady, uniform, flow-field, and the vehicle’s speed, using noisy kinematic measurements during heading change maneuvers, such as circular orbits or 180 degree turns. The three methods proposed included a quadratic curve fitting approach with $(v_{g},\psi)$ data and least-square optimization methods with either $(\dot{x}_{,}\dot{y},\psi)$ or $(v_{g},\psi)$ data. The methods were compared through a Monte Carlo experiment with simulated data sets to illustrate the impact of measurement noise and heading angle change during the maneuver. The comparison included an existing circular curve fitting method from the literature that uses $(\dot{x},\dot{y})$ data. The results indicated that the optimization with $(\dot{x}_{,}\dot{y},\psi)$ led to the lowest mean errors. The methods were also evaluated using a experimental data obtained by an underwater vehicle performing a series of circular orbits during a tidal change. The estimated values had modest agreement with data recorded by a nearby buoy.

### v2

@@ 4 Comparison using Simulated Data
To compare the performance of the four methods described in Secs. 2 and 3, a Monte Carlo experiment was conducted using simulated data. The data was simulated for three different noise settings, $\sigma=\{0.01,0.05,0.10\}$ m/s, and three different heading angle changes, $\Delta\psi=\{2\pi,3\pi/2,\pi\}$. For each combination of $\sigma$ and $\Delta\psi$ a total of 250 unique datasets of $N=100$ noisy samples were generated and transformed to the appropriate form for each algorithm. Each dataset was generated by randomizing the true parameters $(v,w,\theta)$ and drawing them uniformly at random from the interior of the polytope represented by the constraints with $v_{\rm min}=0.5$ and $v_{\rm max}=5$ m/s. For each dataset the circular fit algorithm, optimization with $(\dot{x},\dot{y})$ data, and optimization with $(\dot{v}_{g},\psi)$ data were tested. The quadratic fit approach was evaluated only for the $\Delta\psi=2\pi$ case. The estimate each algorithm produced was compared to the actual value and the absolute error was recorded. The mean of the absolute error, average over 250 trials, is shown in Fig. 6.
As expected, the error for each algorithm increased as the measurement noise is increased and as the range of data is reduced. The quadratic fit method produced the largest errors (e.g., almost 18 deg in mean current direction error for the highest noise setting). The optimization with the $(v_{\rm g},\psi)$ data was not as accurate as the optimization with $(\dot{x},\dot{y},\psi)$ data. This is not surprising since more information is available when the components of the ground velocity are known, rather than the magnitude. The circle fit, which uses $(\dot{x},\dot{y})$ data, in most cases had an intermediate error in comparison to the the two optimization-based methods. The optimization method using $(\dot{x},\dot{y},\psi)$ data produced the most accurate results — for example, in the case of data collected for $\Delta\psi=2\pi$ maneuvers the vehicle and current speed errors were $<$ 0.01 m/s and the current direction error was $<$ 5 degrees.
@@ 5 Comparison using Experimental Data
The four methods described in Secs. 2 and 3 were also evaluated using experimental data obtained by a Bluefin-21 unmanned underwater vehicle operated by the U.S. Naval Research Laboratory. The Bluefin-21 is a 20 ft length, 21 inch diameter vehicle developed by Bluefin Robotics that features a fiber-optic gyroscope inertial navigation system and Doppler velocity log (DVL) navigation suite. The vehicle was deployed on June 22, 2016 near Boston Harbor southeast of Nahant Bay. The vehicle was programmed to execute a series of circular orbits around points of interest to test an onboard sonar system. The ground track of the vehicle is shown in Fig. 7 along with the flow direction estimated by the optimization method with $(\dot{x},\dot{y},\psi)$ data.
@@
The radii of the smaller and larger orbits were 70 and 120 meters, respectively. The data under analysis was truncated to be within 10 meters of these orbits and input into all four estimation algorithms. An example of the dataset along one such orbit is shown in Fig. 8. The parameters estimated by each algorithm were used to superimpose a $v_{\rm g}(\psi)$ curve over the data points. The dataset exhibits outliers that do not conform to the model (e.g., due to transient motions as the vehicle enters and exits each orbit). To address this issue the optimization algorithms are wrapped in a random sample consensus (RANSAC) algorithm outer loop.
The estimated current direction and magnitude was also compared to historical data from an instrumented buoy (identifier: BOS1132) deployed by the National Oceanic and Atmospheric Administration (NOAA) located in Stellwagen Bank about 51 km away directly east from the testing area (15nm NNE of Race Point at a depth of 27ft). The results show good agreement with the ebb and flow tidal directions reported by the buoy. The current magnitude matches the general trend and variations may be due to the spatial separation of the buoy and vehicle leading to different tidal flows closer to shore.
Figure 8: Example experimental data along a single orbit with overlayed $v_{\rm g}(\psi)$ curves for each estimate.
Figure 9: Comparison of estimaties with NOAA buoy data.
@@ 6 Conclusion
Three batch estimation methods were presented that determine the direction and magnitude of a steady, uniform, flow-field, and the vehicle’s speed, using noisy kinematic measurements during heading change maneuvers, such as circular orbits or 180 degree turns. The three methods proposed included a quadratic curve fitting approach with $(v_{\rm g},\psi)$ data and least-square optimization methods with either $(\dot{x}_{,}\dot{y},\psi)$ or $(v_{\rm g},\psi)$ data. The methods were compared through a Monte Carlo experiment with simulated data to illustrate the impact of measurement noise and heading angle change during the maneuver. The comparison included an existing circular curve fitting method from the literature that uses $(\dot{x},\dot{y})$ data. The results indicated that the optimization with $(\dot{x}_{,}\dot{y},\psi)$ led to the lowest mean errors. The methods were also evaluated using a experimental data obtained by an underwater vehicle performing a series of circular orbits during a tidal change. The estimated values had modest agreement with data recorded by a nearby buoy.